# Generative Ghosts: Anticipating Benefits and Risks of AI Afterlives

### v3

@@ Abstract.
As AI systems quickly improve in both breadth and depth of performance, they lend themselves to creating increasingly powerful and realistic agents, including the possibility of agents modeled on specific people. We anticipate that within our lifetimes it may become common practice for people to create a custom AI agent to interact with loved ones and/or the broader world after death; indeed, the past year has seen a boom in startups purporting to offer such services. We call these generative ghosts, since such agents will be capable of generating novel content rather than merely parroting content produced by their creator while living. In this paper, we reflect on the history of technologies for AI afterlives, including current early attempts by individual enthusiasts and by startup companies to create generative ghosts. We then introduce a novel design space detailing potential implementations of generative ghosts, and use this taxonomy to ground discussion of the practical and ethical implications of various approaches to designing generative ghosts, including potential positive and negative impacts on individuals and society. Based on these considerations, we lay out a research agenda for the AI and HCI research communities to better understand the risk/benefit landscape of this novel technology so as to ultimately empower people who wish to create and interact with AI afterlives to do so in a safe and beneficial manner.
@@ 1. Introduction
The past few years have brought incredible growth in the capabilities of generative AI models, particularly large language models (LLMs) such as GPT-4 (OpenAI, 2023a), Palm 2 (Anil et al., 2023), and Llama 2 (Touvron et al., 2023), though there has also been incredible progress in generative AI for the production of images (Ramesh et al., 2022), video (Singer et al., 2022), and audio (Borsos et al., 2023), as well as a new generation of multimodal models (Yang et al., 2023; Google DeepMind, 2023) that combine functionality across several media categories. These models, in turn, have given rise to new types of generative agents (Park et al., 2023), simulacra that can produce believable human behaviors, including capabilities such as memory and planning. While still in their infancy, generative agents and related technologies are likely to increase in fidelity and popularity as underlying model capabilities improve and compute costs drop. For instance, in November 2023 OpenAI released GPTs (OpenAI, 2023b), a no-code interface for people to develop agentic AIs.
As AI models increase the set of human capabilities they can faithfully reproduce (Morris et al., 2023; Bubeck et al., 2023), societal change is inevitable. For instance, experts anticipate that powerful AI systems may profoundly change disparate areas of society such as the labor market (Eloundou et al., 2023), the education system (Kasneci et al., 2023), the pursuit of scientific knowledge (Morris, 2023), and criminal activities (Ferrara, 2023). In this paper, we discuss how advances in AI might change personal and cultural practices around death and dying.
We introduce the concept of generative ghosts, AI agents that represent a deceased person, and discuss why we anticipate such representations may become increasingly popular. We share examples from recent media reports of individuals and startups experimenting with attempting to create generative ghosts, a trend that seems to be accelerating, particularly in East Asia. We introduce a design space of possible instantiations of generative ghosts, and use this taxonomy to structure an analysis of how design choices might lead to practical and ethical concerns as well as potentially beneficial outcomes.
Our contributions include: (1) identifying and characterizing an emerging phenomenon of creating “generative ghosts” to represent the deceased, (2) introducing a taxonomy of design dimensions and analysis of potential benefits and harms that can be used to support future empirical research and motivate fieldwork. By characterizing this emerging trend, highlighting potential risks to be averted, and creating a framework for future investigation, we aim to ensure that any future technical and sociotechnical systems will maximize the potential benefits of “AI afterlives” while minimizing potential risks.
@@ 2.1. Post-Mortem Technology
In HCI, extensive attention has been given to digital legacies, the collection of materials that carry values and meaning that are passed down and/or otherwise continue to represent the deceased after their death. A recent literature review of the field by Doyle and Brubaker (Doyle and Brubaker, 2023) identified four foci of this scholarship: digital identity (i.e., how legacies can continue to represent the deceased in intentional and unintentional ways), engagement with digital legacies (i.e., studies on use and user perceptions), putting to rest (concerned with issues around preservation, disposition, and disposal), and the integration of technology into existing legacy practices (Doyle and Brubaker, 2023).
@@ 2.2. AI Afterlives
A handful of tech-savvy individuals have attempted to create interactive memorializations of loved ones by ingesting their digital (or digitized) content such as emails, journals, videos, photographs, or other autobiographical media. In her graphic novel Artificial: A Love Story, Amy Kurzweil (Kurzweil, 2023) documents how her father, futurist Ray Kurzweil, created a chatbot to embody the memory of his deceased father, Fred Kurzweil, in a form he dubs “Fredbot” (Veltman, 2023). Fredbot interactively responds to questions from his descendants, but only by sharing exact quotes from the materials Fred left behind such as letters digitized by his family. In another well-publicized incident, the engineer Eugenia Kuyda created an app to preserve the memory of her best friend, Roman, who died unexpectedly in an accident, training a neural network on text messages her friend had sent her to create a bot named, suitably, “Roman.” Unlike Fredbot, which was available only to immediate family, the Roman bot was made available on social media and app stores for public interaction, resulting in mixed reactions from friends and family of the deceased (Newton, 2016). While Fredbot and Roman bot represented private citizens, AI has also been used to re-create public figures in various ways. For instance, in November 2023, the surviving members of The Beatles released a new song, “Now and Then,” (Beatles, 2023) using AI to enable the deceased John Lennon to sing along with his living bandmates (Savage, 2023). The musician Laurie Anderson collaborated with the University of Adelaide’s Australian Institute for Machine Learning to create a chatbot based on her longtime partner, deceased musician Lou Reed, which, in addition to conversing, generates new music lyrics in Reed’s style (Andereson, 2024). In early 2024, gun-control activists in the used AI to re-create the voices of victims of gun violence, to create a visceral and poignant plea to U.S. legislators (Luscombe, 2024).
@@
Meanwhile, there are a handful of start-up companies that aim to give people the ability to create their own AI afterlives while they are still alive. Re;memory (rememory.deepbrain.io/en) is an offering from DeepBrain AI that professes to create an interactive virtual representation after a seven-hour filming and interview session. They specifically advertise this service as a way that one can proactively create a rich memory that friends and family can engage with after one dies. HereAfter (https://www.hereafter.ai) provides an app that interviews a user with the goal of proactively creating a posthumous digital representation. Friends and relatives can interact with a chatbot-based representation of their loved one who can provide photos and voice recordings of memories and life events. Early attempts at generative ghost technologies seem to already be entering the mainstream in East Asia, where ”communicating” (via non-technological means) with deceased ancestors is already a cultural norm for many (Yang, 2024a). Companies offering ”digital immortality” are increasingly popular in China, serving thousands of customers (Yang, 2024b; Feng, 2024; Yang, 2024a). In South Korea, 19 million people have streamed an emotional video of a bereaved mother, Jang Ji-Sun, interacting with a virtual reality representation of her deceased young daughter, Na-yeon, custom-created for her by a media company (Chesler, 2022).
@@
In contrast with Re;memory and HereAfter, which aim to empower users to preserve their memories before their deaths, some start-ups purport to use AI to provide an experience more akin to resurrection or reanimation (i.e., without the consent of the deceased party). For instance, Character.AI (https://beta.character.ai/) uses LLMs to simulate chatting with deceased public figures for entertainment or education (e.g., a bot representing William Shakespeare), and the Musée D’Orsay in Paris recently contracted with a start-up company to develop a chatbot meant to represent the artist Vincent van Gogh by ingesting his writings (Small, 2023). Beyond public figures, Project December (projectdecember.net/) is a venture that allows end-users to upload content from a deceased loved one (Fagone, 2021) in order to “simulate the dead” via a customized chatbot. The Augmented Eternity project at MIT is an academic endeavor that aims to allow people to create digital representations of themselves with the express purpose of being able to agentically represent them after death to members of their social network (Rahnama et al., 2021).
@@
Because of their third-party origins, griefbots raise controversial ethical issues regarding privacy and consent (Grandinetti et al., 2020) which must be considered from multiple perspectives. For example, Hollanek and Nowaczyk‑Basińska (Hollanek and Nowaczyk-Basińska, 2024) emphasize the importance of considering the complex relationships among three classes of stakeholders for such bots: data donors, data recipients, and service interactants. These align with our own thinking, both in our design framework (see first- vs. third-party generative ghosts) and in our discussion of risks and benefits. However, we use the terms deceased and bereaved to emphasize interpersonal and relational aspects of clones, while also considering the impacts on society at large.
Generative ghosts can be considered an extension of the concept of a griefbot, but not all griefbots would be considered generative ghosts. Generative ghosts go beyond the traditional definition of griefbots in several ways: they always include the ability to generate novel content in-character (whereas many griefbots only regurgitate content from a training corpus), potentially evolve over time, and possess agentic (Park et al., 2023; Shavit et al., 2023) capabilities such as the ability to participate in the economy or perform other complex tasks with limited oversight. Notably, our framework for generative ghosts envisions the possibility of first-party (rather than only third-party) creation as part of an end-of-life planning process. Also, generative ghosts may be designed to represent non-human entities (such as pets), and in many cases might exist pre-mortem as a generative clone. Finally, in contrast to griefbots, generative ghosts may be created for reasons other than to support the bereaved in their grieving process. The following section introduces more details about the design space of generative ghosts.
@@ 3. Generative Ghosts: A Design Space
We introduce the notion of a generative ghost, an agentic, AI-powered representation of a deceased individual. Generative ghosts differ from other technology-mediated representations of the deceased (such as memorialized social media accounts) in that they provide more than static or interactive access to content about the deceased produced during their lifetime (e.g., text, imagery, audio, video, or other media). Rather, they additionally are capable of generating novel content and interactions based upon the deceased’s data, potentially mediated by additional metadata such as personality questionnaires (NORC at the University of Chicago, nd), prompts (Zamfirescu-Pereira et al., 2023), constitutions (Bai et al., 2022), or other rules set forth by the generative ghost’s creator. In addition to producing novel content, generative ghosts may also possess agentic (Park et al., 2023; Shavit et al., 2023) capabilities (i.e., the ability to autonomously execute complex sequences of actions with side effects in either the digital or physical realms).
@@
The provenance of a generative ghost refers to who created it (or authorized its creation). A first-party generative ghost is one created by the individual being represented (i.e., perhaps via end-of-life planning or legacy crafting practices). In contrast, a third-party generative ghost is not created by the representee. Third-party generative ghosts might be created by parties with a direct personal connection to the deceased (e.g., family, friends), by parties with a fiscal connection to the deceased (e.g., employers, estates), or by unconnected parties. Additionally, third-party ghosts may be authorized or unauthorized by the representee. Authorized third-party generative ghosts, for example, might be created with consent (e.g., via the deceased’s will). Meanwhile, unauthorized ghosts created by unconnected third parties are most likely in the case of public figures such as historical figures or contemporary celebrities; indeed, we already see early attempts at creating third-party generative ghosts of historical figures for entertainment and educational purposes by companies such as character.ai (Metz, 2023) and nonprofits such as Khan Academy (Grant, 2023).
@@
The deployment timeline of a generative ghost indicates whether the system’s initial deployment was pre-mortem or post-mortem. While some generative ghosts will be deployed post-mortem with the explicit purpose of memorializing the dead, many generative ghosts will likely start off as agents that serve some pre-mortem functionality (e.g., responding to personal emails or accomplishing other tasks on behalf of the representee, in character as if the representee had executed the task themself) – what we have termed generative clones – only to transition to ghosts after the death of their representee. Pre-mortem deployments might be desirable as they would allow an individual the opportunity to tune the behavior and/or capabilities of a first-party generative ghost to their liking, perhaps as part of an end-of-life planning process. Likewise, we imagine that generative clones might be designed with mortality in mind, potentially with attendant modifications to their behavior and/or capabilities once they become ghosts. Note that the ethical implications of pre-mortem generative clones (also called “digital twins,” “AI clones,” or “mimetic models”) have been discussed in related works such as (McIlroy-Young et al., 2022; Lee et al., 2023); our contribution focuses on identifying potential risks, benefits, and a research agenda around post-mortem clones, i.e., generative ghosts.
@@
The anthropomorphism paradigm of a generative ghost refers to a subtle, but potentially impactful, interface metaphor choice: whether a generative ghost presents itself as a reincarnation of the deceased individual or as a representation of that individual. Generative ghosts employing a reincarnation metaphor would engage with end-users as if they were the deceased individual. In contrast, generative ghosts employing a representation metaphor would engage with end-users as if they are a representation of the deceased, but are not the deceased themselves. Design choices that might impact perceptions of whether a generative ghost is a reincarnation or a representation include whether a generative ghost uses first-person pronouns, whether it uses the present or past tense when discussing its representee, whether it uses the name of its representee or a different nomenclature (e.g., the use of “Fredbot” in (Kurzweil, 2023)), whether it is allowed to make statements that assert it is alive, possesses a soul, etc.
@@
While the previous example speaks to whether a ghost is designed to evolve by simulating aging, ghosts will also evolve (albeit in less predictable ways) should new information be added to their models (e.g., in the case that additional data about the deceased was found). Moreover, the informational context beyond data about the representee (e.g., from world news to the personal information of those who might interact with the ghost) can be static or be allowed to evolve. One can imagine scenarios in which a ghost might be designed to incorporate (or not) new information, including personal information (such as a wedding or birth of a child) or world events (such as a future election).

### v4

@@ Abstract.
As AI systems quickly improve in both breadth and depth of performance, they lend themselves to creating increasingly powerful and realistic agents, including the possibility of agents modeled on specific people. We anticipate that within our lifetimes it may become common practice for people to create custom AI agents to interact with loved ones and/or the broader world after death; indeed, the past year has seen a boom in startups purporting to offer such services. We call these generative ghosts, since such agents will be capable of generating novel content rather than merely parroting content produced by their creator while living. In this paper, we reflect on the history of technologies for AI afterlives, including current early attempts by individual enthusiasts and startup companies to create generative ghosts. We then introduce a novel design space detailing potential implementations of generative ghosts, and use this analytic framework to ground discussion of the practical and ethical implications of various approaches to designing generative ghosts, including potential positive and negative impacts on individuals and society. Based on these considerations, we lay out a research agenda for the AI and HCI research communities to better understand the risk/benefit landscape of this novel technology so as to ultimately empower people who wish to create and interact with AI afterlives to do so in a beneficial manner.
@@ 1. Introduction
The past few years have brought incredible growth in the capabilities of generative AI models, particularly large language models (LLMs) such as GPT-4 (OpenAI, 2023a), Palm 2 (Anil et al., 2023), and Llama 2 (Touvron et al., 2023), though there has also been incredible progress in generative AI for the production of images (Ramesh et al., 2022), video (Singer et al., 2022), and audio (Borsos et al., 2023), as well as a new generation of multimodal models (Yang et al., 2023; Google DeepMind, 2023) that combine functionality across several media categories. These models, in turn, have given rise to new types of generative agents (Park et al., 2023), simulacra that can produce believable human behaviors, including capabilities such as memory and planning. While still in their infancy, generative agents and related technologies are likely to increase in fidelity and popularity as underlying model capabilities improve, compute costs drop, and the required technical expertise decreases. For instance, in November 2023 OpenAI released GPTs (OpenAI, 2023b), a no-code interface for people to develop agentic AIs. As AI models increase the set of human capabilities they can faithfully reproduce (Morris et al., 2023; Bubeck et al., 2023), societal change is inevitable. For instance, experts anticipate that powerful AI systems may profoundly change disparate areas of society such as the labor market (Eloundou et al., 2023), the education system (Kasneci et al., 2023), the pursuit of scientific knowledge (Morris, 2023), and criminal activities (Ferrara, 2023). In this paper, we discuss how advances in AI might change personal and cultural practices around death and dying as well.
We introduce the concept of generative ghosts [^1] We use the term ”ghosts” to encapsulate both the memorializing practices we imagine these agents may fulfill and the unintentional persistent presences we anticipate such AI agents will bring., AI agents that represent a deceased person, and discuss why we anticipate such representations may become increasingly common. We share examples from recent media reports of individuals and startups experimenting with the creation of generative ghosts, a trend that seems to be accelerating, particularly in East Asia. We introduce a design space to describe possible instantiations of generative ghosts and use this analytic framework to explore how design choices might lead to practical and ethical concerns as well as potentially beneficial outcomes.
Our contributions include: (1) identifying and characterizing an emerging phenomenon of creating “generative ghosts” to represent the deceased, (2) introducing a design space of dimensions and analysis of potential benefits and harms that can be used to support future empirical research and motivate fieldwork. By characterizing this emerging trend, highlighting potential risks, and creating a framework for future investigation, we aim to ensure that future technical and sociotechnical systems will maximize the potential benefits of “AI afterlives” while minimizing potential risks.
@@ 2.1. Post-Mortem Technology
In HCI, extensive attention has been given to digital legacies, the collection of materials that carry values and meaning that are passed down and/or otherwise continue to represent the deceased after their death. A recent literature review of the field by Doyle and Brubaker (Doyle and Brubaker, 2023) identified four foci of this scholarship: digital identity (i.e., how legacies can continue to represent the deceased in intentional and unintentional ways), engagement with digital legacies (i.e., studies on use and user perceptions), putting to rest (i.e., concerned with issues around preservation, disposition, and disposal), and the integration of technology into existing legacy practices (Doyle and Brubaker, 2023).
@@ 2.2. AI Afterlives
A handful of tech-savvy individuals have attempted to create interactive memorializations of loved ones by ingesting their digital (or digitized) content such as emails, journals, videos, photographs, or other autobiographical media. In her graphic novel Artificial: A Love Story, Amy Kurzweil (Kurzweil, 2023) documents how her father, futurist Ray Kurzweil, created a chatbot to embody the memory of his deceased father, Fred Kurzweil, in a form he dubs “Fredbot” (Veltman, 2023). Fredbot interactively responds to questions from his descendants, but only by sharing exact quotes from the materials Fred left behind such as letters digitized by his family. In another well-publicized incident, the engineer Eugenia Kuyda created an app to preserve the memory of her best friend, Roman, who died unexpectedly in an accident, training a neural network on text messages her friend had sent her to create a bot named, suitably, “Roman.” Unlike Fredbot, which was available only to immediate family, the Roman bot was made available on social media and app stores for public interaction, resulting in mixed reactions from friends and family of the deceased (Newton, 2016). While Fredbot and Roman bot represented private citizens, AI has also been used to re-create public figures in various ways. For instance, in November 2023, the surviving members of The Beatles released a new song, “Now and Then,” (Beatles, 2023) using AI to enable the deceased John Lennon to sing along with his living bandmates (Savage, 2023). The musician Laurie Anderson collaborated with the University of Adelaide’s Australian Institute for Machine Learning to create a chatbot based on her longtime partner, deceased musician Lou Reed, which, in addition to conversing, generates new music lyrics in Reed’s style (Andereson, 2024). In early 2024, gun-control activists in the United States used AI to re-create the voices of victims of gun violence, to create a visceral and poignant plea to U.S. legislators (Luscombe, 2024).
@@
Meanwhile, there are a handful of start-up companies that aim to give people the ability to create their own AI afterlives while they are still alive. Re;memory222rememory.deepbrain.io/en is an offering from DeepBrain AI that professes to create an interactive virtual representation after a seven-hour filming and interview session. They specifically advertise this service as a way that one can proactively create a rich memory that friends and family can engage with after one dies. HereAfter [^3] https://www.hereafter.ai provides an app that interviews a user with the goal of proactively creating a posthumous digital representation. Friends and relatives can interact with a chatbot-based representation of their loved one who can provide photos and voice recordings of memories and life events. Early attempts at generative ghost technologies seem to already be entering the mainstream in East Asia, where ”communicating” (via non-technological means) with deceased ancestors is already a cultural norm for many (Yang, 2024a). Companies offering ”digital immortality” are increasingly popular in China, serving thousands of customers (Yang, 2024b; Feng, 2024; Yang, 2024a). In South Korea, 19 million people have streamed an emotional video of a bereaved mother, Jang Ji-Sun, interacting with a virtual reality representation of her deceased young daughter, Na-yeon, custom-created for her by a media company (Chesler, 2022).
@@
In contrast with Re;memory and HereAfter, which aim to empower users to preserve their memories before their deaths, some start-ups purport to use AI to provide an experience more akin to resurrection or reanimation (i.e., without the direct involvement of the representee, and often after their death). For instance, Character.AI [^4] https://beta.character.ai/ uses LLMs to simulate chatting with deceased public figures for entertainment or education (e.g., a bot representing William Shakespeare), and the Musée D’Orsay in Paris recently contracted with a start-up company to develop a chatbot meant to represent the artist Vincent van Gogh by ingesting his writings (Small, 2023). Beyond public figures, Project December555projectdecember.net/ is a venture that allows end-users to upload content from a deceased loved one (Fagone, 2021) in order to “simulate the dead” via a customized chatbot. The Augmented Eternity project at MIT is an academic endeavor that aims to allow people to create digital representations of themselves with the express purpose of being able to agentically represent them after death to members of their social network (Rahnama et al., 2021).
@@
Because of their third-party origins, griefbots raise controversial ethical issues regarding privacy and consent (Grandinetti et al., 2020) which must be considered from multiple perspectives. For example, Hollanek and Nowaczyk‑Basińska (Hollanek and Nowaczyk-Basińska, 2024) emphasize the importance of considering the complex relationships among three classes of stakeholders for such bots: data donors, data recipients, and service interactants. The complexity they raise aligns with our own thinking, both in our design framework (see first- vs. third-party generative ghosts) and in our discussion of risks and benefits. However, we use the terms deceased and bereaved to emphasize interpersonal and relational aspects of clones, while also considering the impacts on society at large.
Generative ghosts can be considered an extension of the concept of a griefbot, but not all griefbots would be considered generative ghosts. Generative ghosts go beyond the traditional definition of griefbots in several ways. Generative ghosts always include the ability to generate novel content in-character (whereas many griefbots only regurgitate content from a training corpus) and potentially evolve over time. For instance, a user should be able to ask a generative ghost questions about current events and obtain responses that would be “in character” for the deceased even though these events occurred after the end of the representee’s life. Additionally, generative ghosts extend the notion of griefbots by possessing agentic (Park et al., 2023; Shavit et al., 2023) capabilities such as the ability to participate in the economy or perform other complex tasks with limited oversight (in AI parlance, generative ghosts can engage in tool use (Bubeck et al., 2023)). Notably, our framework for generative ghosts envisions the possibility of first-party (rather than only third-party) creation as part of an end-of-life planning process. Another key expansion to the notion of griefbots is that generative ghosts might exist pre-mortem as a generative clone (an AI agent representing a living person), which then transitions into a generative ghost upon the representee’s death. Finally, in contrast to griefbots, generative ghosts may be created for reasons other than to support the bereaved in their grieving process. The following section introduces more details about the design space of generative ghosts.
@@ 3. Generative Ghosts: A Design Space
We introduce the notion of a generative ghost, an agentic, AI-powered representation of a deceased individual. Drawing on methods from design research (e.g., (Yildirim et al., 2023)), we used an analytical approach to generate a design space. Our analysis started with a growing set of individual attempts to create generative ghosts as reported in the news media and descriptions of start-up companies purporting to offer related services (see Section 2). Drawing on our respective expertise as scholars in technical human-AI interaction (first author) and in thanatosensitive design (second author) we considered those systems and services in the context of related work in our respective areas, speculating on possible and likely futures. As such, our proposed framework is the byproduct of analytic and conceptual explorations drawing on related work and systems in the wild to date. We share this framework with the hopes that it will guide future work while also knowing that future design and empirical work will likely result in updates to our proposed framework.
Generative ghosts differ from other technology-mediated representations of the deceased (such as memorialized social media accounts) in that they provide more than static or interactive access to content about the deceased produced during their lifetime (e.g., text, imagery, audio, video, or other media). Rather, they additionally are capable of generating novel content and interactions based upon the deceased’s data, potentially mediated by additional metadata such as personality questionnaires (NORC at the University of Chicago, nd), prompts (Zamfirescu-Pereira et al., 2023), constitutions (Bai et al., 2022), or other rules set forth by the generative ghost’s creator. In addition to producing novel content, generative ghosts may also possess agentic (Park et al., 2023; Shavit et al., 2023) capabilities (i.e., the ability to autonomously execute complex sequences of actions with side effects in either the digital or physical realms).
@@
The provenance of a generative ghost refers to who created it (or authorized its creation). A first-party generative ghost is one created by the individual being represented (i.e., perhaps via end-of-life planning or legacy crafting practices). First-party ghosts may also be created by the individual, but not explicitly for legacy purposes (akin to how social media profiles may become memorials after someone dies (Doyle and Brubaker, 2023)). In contrast, a third-party generative ghost is not created by the representee. Third-party generative ghosts might be created by parties with a direct personal connection to the deceased (e.g., family, friends), by parties with a fiscal connection to the deceased (e.g., employers, estates), or by unconnected parties. Additionally, third-party ghosts may be authorized or unauthorized by the representee. Authorized third-party generative ghosts, for example, might be created with consent (e.g., via the deceased’s will). Meanwhile, unauthorized ghosts created by unconnected third parties are most likely in the case of public figures such as historical figures or contemporary celebrities; indeed, we already see early attempts at creating third-party generative ghosts of historical figures for entertainment and educational purposes by companies such as character.ai (Metz, 2023) and nonprofits such as Khan Academy (Grant, 2023).
@@
Closely related to provenance, the deployment timeline of a generative ghost indicates whether the system’s initial deployment was pre-mortem or post-mortem. While some generative ghosts will be deployed post-mortem with the explicit purpose of memorializing the dead, many generative ghosts will likely start off as agents that serve some pre-mortem functionality (e.g., a generative clone responding to personal emails or accomplishing other tasks on behalf of the representee, in character as if the representee had executed the task themself) only to transition to ghosts after the death of their representee. Similar to digital legacy planning (Doyle and Brubaker, 2024, 2023), pre-mortem deployments might be desirable as they would allow an individual the opportunity to tune the behavior and/or capabilities of a first-party generative ghost to their liking, perhaps as part of an end-of-life planning process. We also expect that generative clones would benefit from being designed with mortality in mind, in line with calls for thanatosensitivity in other technology (Massimi and Charise, 2009). The design of clones should likely include modifications to their behavior and/or capabilities once they become ghosts. [^6] Note that the ethical implications of pre-mortem generative clones (also called “digital twins,” “AI clones,” or “mimetic models”) have been discussed in related works such as (McIlroy-Young et al., 2022; Lee et al., 2023); our contribution focuses on identifying potential risks, benefits, and a research agenda around post-mortem clones, i.e., generative ghosts.
@@
The anthropomorphism paradigm of a generative ghost refers to a subtle, but potentially impactful, interface metaphor choice: whether a generative ghost presents itself as a reincarnation of the deceased individual or as a representation of that individual. Generative ghosts employing a reincarnation metaphor would engage with end-users as if they were the deceased individual. In contrast, generative ghosts employing a representation metaphor would engage with end-users as if they are a representation of the deceased, but are not the deceased themselves. We see similar distinctions in the context of social media where HCI literature and popular press have documented scenarios where platforms animate users even after their deaths (e.g., through algorithmically generated notifications (Jiang and Brubaker, 2018)) and the need to be clear about whether content is the result of a user’s actions or the platform’s (Brubaker et al., 2013). In the context of generative ghosts, some design choices that might impact perceptions of reincarnation vs. representation include whether a generative ghost uses first-person pronouns, whether it uses the present or past tense when discussing its representee, whether it uses the name of its representee or a different nomenclature (e.g., the use of “Fredbot” in (Kurzweil, 2023)), and whether it is allowed to make statements that assert it is alive, possesses a soul, etc.
@@
While the previous example speaks to whether a ghost is designed to evolve by simulating aging, a ghost could also evolve (albeit in less predictable ways) if new information about the deceased is added to the model (e.g., in the case that additional data about the deceased was found). Moreover, the informational context beyond data about the representee (e.g., from world news to the personal information of those who might interact with the ghost) can be static or be allowed to evolve. One can imagine scenarios in which a ghost might be designed to incorporate (or not) new information, including personal information (such as a wedding or birth of a child) or world events (such as a future election). [^7] In 2013, a now-defunct service called LivesOn aimed to approximate a rudimentary version of this concept. The service promised to animate a deceased user’s Twitter profile with tweets written in their likeness, commenting on the latest news headlines (Gayle, 2013).